OpenStudy (ttop0816):

Evaluate AB/CD for A = 5, B = -2, C = 4 and D = -6. 5/12 -5/12 -3/2 3/2

OpenStudy (anonymous):

AB / CD (A = 5, B = -2, C = 4, D = -6) now just sub (5 * -2) / (4 * -6) can you finish ?

OpenStudy (anonymous):

Its nothing but = A*B/C*D

OpenStudy (ttop0816):

@kelliegirl33 5/12!

OpenStudy (anonymous):

Yes you are exactly correct,,,,,@ttop0816

OpenStudy (anonymous):

-10/-24 = 10/24 = 5/12....you got it

OpenStudy (ttop0816):

could you also help me with this?! Select the mathematical terms that mean "opposite of." reciprocal negative additive inverse multiplicative inverse minus absolute value

OpenStudy (anonymous):

I believe it is additive inverse